OpenStudy (anonymous):

USe Squeeze Theorem to determine lim x->0 (x+1) cos( ln x^2) / √x^2+2

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ (x+1)\cos (\ln x ^{2)} }{ \sqrt{x ^{2} +2} }\]

OpenStudy (sidsiddhartha):

start with \[-1 \le \cos(\ln x^2) \le1\]

OpenStudy (sidsiddhartha):

so \[\frac{ (x+1) }{ \sqrt{x^2+2 }} \le \frac{ (x+1) }{ \sqrt{x+2} }\cos(\ln x^2) \le -\frac{ (x+1) }{ \sqrt{x+2}}\] getting this?

OpenStudy (sidsiddhartha):

now according to squeeze theorem evaluate \[\lim_{x \rightarrow 0}\frac{ (x+1) }{ \sqrt{x+2} }\] can u do that?

OpenStudy (sidsiddhartha):

ok just simply evaluate the limit by setting x=0 \[\lim_{x \rightarrow 0}\frac{ (x+1) }{ \sqrt{x+2} }=\frac{ 0+1 }{ \sqrt{0+2} }=\frac{ 1 }{ \sqrt{2} }\] and for the left hand side \[\lim_{x \rightarrow 0}\frac{ -(x+1) }{ \sqrt{x+2} }=\frac{ -1 }{ \sqrt{2} }\] so u can see \[Left.hand.limit \neq Right.hand.limit\] so \[LIMIT~DOES~NOT~EXISTS\]

OpenStudy (sidsiddhartha):

@jibirajeev

OpenStudy (anonymous):

let me try. in squeeze theorem we have to get two points one from left and right and our answer must be inbetween. am i right?

OpenStudy (anonymous):

Thanks.. got it..:)

OpenStudy (sidsiddhartha):

yeah thats how we proceed YW!!!!

OpenStudy (anonymous):

Thanks... i do understand what you done for me here. if you dont mind can you explain one question with limit does exist.